OpenStudy (anonymous):

One angle of a right triangle measures 55 . Which could be the measure of one of the other angles? 35 45 55 125

OpenStudy (anonymous):

35

OpenStudy (anonymous):

35

OpenStudy (lalaly):

180-(90+55) =35

OpenStudy (anonymous):

55 + 90+x = 180 x = 180-(55+90) x = 180-145 x= 35

OpenStudy (anonymous):

well since it is a right triangle one of the other sides has to be 90 so u have one side with 55 and one side with 90. if u add those 2gether, u get 145. Each triangle adds up to 180, so if u subtract 145 from 180 u get 35, therefore making the last angle of the triangle 35.